### v1

# ROTI-GCV: Generalized Cross-Validation for right-ROTationally Invariant Data

Abstract
Two key tasks in high-dimensional regularized regression are tuning the regularization strength for good predictions and estimating the out-of-sample risk. It is known that the standard approach — $k$-fold cross-validation — is inconsistent in modern high-dimensional settings. While leave-one-out and generalized cross-validation remain consistent in some high-dimensional cases, they become inconsistent when samples are dependent or contain heavy-tailed covariates. To model structured sample dependence and heavy tails, we use right-rotationally invariant covariate distributions - a crucial concept from compressed sensing. In the common modern proportional asymptotics regime where the number of features and samples grow comparably, we introduce a new framework, ROTI-GCV, for reliably performing cross-validation. Along the way, we propose new estimators for the signal-to-noise ratio and noise variance under these challenging conditions. We conduct extensive experiments that demonstrate the power of our approach and its superiority over existing methods.
1 Introduction
Tuning the regularization strength in penalized regression estimators is a fundamental technique for addressing overfitting in machine learning and data science, particularly important for modern high-dimensional datasets. In this work, we propose a new cross-validation technique, ROTI-GCV, for determining one of the most fundamental hyperparameters, the ridge regularization parameter $\lambda$, under a different model on the design $\mathbf{X}$. The classical approach for tuning $\lambda$ is $k$-fold cross-validation [20], where the training data is divided into $k$ equally sized chunks (folds). A model is then repeatedly trained on all but one fold and validated on the remaining fold. The value of $\lambda$ achieving the best average validation performance is then selected. However, in high-dimensional settings where the dataset dimension $p$ scales proportionally with dataset size $n$, using a finite $k$ results in an inconsistent estimate of out-of-sample error [32, 43]. On the other hand, leave-one-out cross-validation (LOOCV, corresponding to $k=n$) and the lesser-known generalized cross-validation (GCV) remain consistent in some high-dimensional settings [19, 3, 31].
The theoretical consistency of LOOCV and GCV in high dimensions has thus far been based on the assumption that the samples are independent and identically distributed (i.i.d.) draws from some fixed distribution of sufficient regularity, such as having light enough tails [20, 3, 31]. These assumptions, however, do not always hold in practice. We show in this study that in settings where there is dependence across observations or that the data is heavy-tailed, LOOCV and GCV become biased. To address the limitations of i.i.d. assumptions, we propose modeling designs using an alternative random matrix ensemble. Instead of assuming the rows of our design to be i.i.d. draws from a fixed distribution, we instead require that the singular value decomposition of the design $\mathbf{X}$ is right-rotationally invariant (see Definition 1).
Right-rotationally invariant designs form a rich class, allowing for tractable analysis while also capturing structured forms of dependence among datapoints. These designs have recently received much attention as alternatives to fully i.i.d. designs for theoretical analysis in a multitude of works [38, 29, 33, 39, 7, 11, 14, 40, 41, 42, 15, 28, 26, 46, 27, 13, 12]. Recent works [8, 44] also show that, under certain conditions, the behavior of high dimensional matrices is largely dictated by their spectrum, and in fact will “act” similarly to a right-rotationally invariant matrices with the same spectrum.
In this work, we introduce a new framework inspired by generalized cross-validation, which we call ROTI-GCV, that is provably consistent for the out-of-sample risk when the data is right-rotationally invariant. We then provide simulations demonstrating that the new formulation outperforms traditional GCV and LOOCV in scenarios where observations are correlated and exhibit heavy tails.
2 Preliminaries and Setup
2.1 Right-rotationally invariant designs
Definition 1 (Right-rotationally invariant design).
A random design matrix $\mathbf{X}\in\mathbb{R}^{n\times p}$ is right-rotationally invariant if for any $\mathbf{O}\in\mathbb{O}(p)$, one has $\mathbf{XO}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}$, where $\mathbb{O}(p)$ denotes the group of $p\times p$ orthogonal matrices. Equivalently, let $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{D}\mathbf{O}$ be the singular value decomposition of $\mathbf{X}$. Then $\mathbf{X}$ is right-rotationally invariant if and only if $\mathbf{O}$ is independent of $(\mathbf{Q},\mathbf{D})$ and $\mathbf{O}$ drawn from the Haar measure on $\mathbb{O}(p)$ that is, it is uniformly distributed on $\mathbb{O}(p)$.
The setting where each entry of $\mathbf{X}$ is an i.i.d. standard Gaussian is the simplest example of a right-rotationally invariant design. Other examples of right-rotationally invariant designs include the following:
  • Equicorrelated data: $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$. This can model situations where the training dataset a populations sharing a latent factor (e.g. location), causing the samples to have correlated covariates.
  • Autocorrelated data: The rows of $\mathbf{X}$ can be drawn from a time series, where $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\boldsymbol{\epsilon}_{i}$, with $\boldsymbol{\epsilon}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$.
  • $t$-distributed data: The rows of $\mathbf{X}$ can be drawn from a multivariate $t$ distribution with as few as $3$ degrees of freedom, capturing heavy-tailed data such as financial returns [16].
  • products of Gaussian matrices: $\mathbf{X}=\mathbf{X}_{1}\mathbf{X}_{2}\cdot\cdots\cdot\mathbf{X}_{k}$, where $\mathbf{X}_{1}$ has $n$ rows and $\mathbf{X}_{k}$ has $p$ columns, while the remaining dimensions are arbitrary. See [17, 18] for connections to linear neural networks.
  • spiked matrices: $\mathbf{X}=\lambda\mathbf{V}\mathbf{W}^{\top}+\mathbf{G}$, where $\mathbf{V}\in\mathbb{R}^{n\times r}$ and $\mathbf{W}\in\mathbb{R}^{p\times r}$ are the first $r$ columns of two Haar matrices.
See [27, Figure 1] for more such examples.
In essence, right-rotationally invariant ensembles allow us enormous flexibility—they can simultaneously capture dependent rows and heavy-tails by allowing a rather general class of singular value distributions. The spectrum of right-rotationally invariant designs can essentially be arbitrary, and allowing for significant deviations from prior i.i.d. designs used to study the behavior of LOOCV and GCV in high dimensions. Furthermore, their universality class is extremely broad, as established in [8], capturing even challenging deterministic designs arising in compressed sensing. At the same time, the fact that $\mathbf{O}$ is Haar does add some restrictions - in particular, the designs we consider are now necessarily isotropic.
Lemma 1.
Let $\mathbf{X}=\mathbf{Q}^{\top}\mathbf{DO}$ be right-rotationally invariant. Then $\mathbb{E}[\mathbf{X}^{\top}\mathbf{X}]=\mathsf{Tr}(\mathbb{E}[\mathbf{D}^{\top}\mathbf{D}])/p\cdot\mathbf{I}_{p}$.
In prior analyses, the spectrum of $\boldsymbol{\Sigma}$ defines the problem geometry. We sacrifice this degree of freedom for the ability to set the spectrum $\mathbf{D}$ to capture distributions beyond those of i.i.d. rows.
To simplify some statements, we introduce some notation. Let $m(\mu,z)$ denote the Stieltjes transform of a measure $\mu$ on $\mathbb{R}$. In our analysis, we will mostly care when $\mu$ is the empirical distribution of the entries of $\mathbf{D}^{\top}\mathbf{D}$ (which are the same as those of $\mathbf{X}^{\top}\mathbf{X}$). Take $\mu_{\mathbf{D}^{\top}\mathbf{D}}$ to denote the empirical measure of the diagonal entries of $\mathbf{D}^{\top}\mathbf{D}$, and let $m_{\mathbf{D}}(z)=m(\mu_{\mathbf{D}^{\top}\mathbf{D}},z)$. We also consider the companion Stieltjes transform, which is the Stieltjes transform of the empirical distribution of $\mathbf{D}\mathbf{D}^{\top}$. We will denote this as $v_{\mathbf{D}}(z)=m(\mu_{\mathbf{D}\mathbf{D}^{\top}},z)$.
2.2 Problem Setup
We study high-dimensional linear regression with right-rotationally invariant design. Concretely, one observes $(\mathbf{X},\mathbf{y})$, such that
$\mathbf{y}=\mathbf{X}\boldsymbol{\beta}+\boldsymbol{\epsilon},$ (1)
with $\mathbf{y}\in\mathbb{R}^{n}$, $\mathbf{X}\in\mathbb{R}^{n\times p}$, and $\boldsymbol{\epsilon},\boldsymbol{\beta}\in\mathbb{R}^{p}$. Hence $n$ is the sample size, $p$ is both the data dimension and the parameter count, $\boldsymbol{\beta}$ is the signal, and $\boldsymbol{\epsilon}$ is a component-wise independent noise vector, where each entry has mean 00 and variance $\sigma^{2}$. We model $\mathbf{X}$ as being right-rotationally invariant, $\boldsymbol{\beta}$ as being fixed or random independent of $\mathbf{X}$, and $\boldsymbol{\epsilon}$ independent of everything else.
The problem we consider will be the expected out-of-sample risk on an independent draw from a new, potentially different right-rotationally invariant design, $\widetilde{\mathbf{X}}=\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$, with $\widetilde{\mathbf{X}}\in\mathbb{R}^{n^{\prime}\times p}$. This corresponds, for example, to the setting of learning a model on one interdependent population (such as medical data from some population), and using the model on a separate interdependent population. The explicit risk we examine, following [31], is
$R_{\mathbf{X},\mathbf{y}}\left(\hat{\boldsymbol{\beta}}(\mathbf{X},\mathbf{y}),\boldsymbol{\beta}\right)=\frac{1}{n^{\prime}}\mathbb{E}\left[\left.\left\|\widetilde{\mathbf{X}}\hat{\boldsymbol{\beta}}-\widetilde{\mathbf{X}}\boldsymbol{\beta}\right\|^{2}\right|\mathbf{X},\mathbf{y}\right].$ (2)
This setting is comparable to the isotropic setting in [19].
We study the risk of the ridge estimator
$\hat{\boldsymbol{\beta}}_{\lambda}=\operatorname*{arg\;min}_{\mathbf{b}\in\mathbb{R}^{p}}\left\{\|\mathbf{y}-\mathbf{X}\mathbf{b}\|_{2}^{2}+\lambda\|\mathbf{b}\|^{2}_{2}\right\}$ (3)
which has closed form $\hat{\boldsymbol{\beta}}=(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{y}$.
2.2.1 Scaling and Intuition
Lemma 1 shows that $\mathbb{E}[\mathbf{X}^{\top}\mathbf{X}]=\mathcal{O}(1)\cdot\mathbf{I}$, while in the i.i.d. setting with identity covariance, one would expect $\mathbb{E}[\mathbf{X}^{\top}\mathbf{X}/n]=\mathbf{I}$. Thus, $\mathbf{X}$ in our setting is akin to the normalized matrix $\mathbf{X}/\sqrt{n}$ in the i.i.d. setting. Furthermore, $\mathbf{D}^{\top}\mathbf{D}$ contains the eigenvalues of $\mathbf{X}^{\top}\mathbf{X}$, and this scaling ensures that $\|\mathbf{D}^{\top}\mathbf{D}\|_{\mathsf{op}}=\mathcal{O}(1)$ in the setting where entries are i.i.d. draws from some fixed distribution. The consequence of this scaling is that in order to maintain the signal to noise ratio, we must have $\boldsymbol{\beta}$ having norm on the order of $\sqrt{n}$.
3 Generalized Cross-Validation
Before stating our results, we note the running assumptions used in what follows. We consider a sequence of right-rotationally invariant designs $\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{n},\dots$ where each $\mathbf{X}_{i}\sim\mathbf{Q}_{i}^{\top}\mathbf{D}_{i}\mathbf{O}_{i}$, and a sequence of signal vectors $\boldsymbol{\beta}_{1},\boldsymbol{\beta}_{2},\dots,\boldsymbol{\beta}_{n},\dots$ such that for all $n$, $\boldsymbol{\beta}_{n}\in\mathbb{R}^{p}$ and $\|\boldsymbol{\beta}_{n}\|/\sqrt{n}=r$. We generate $\mathbf{y}_{i}$ according to (1). In the sequel, we drop the dependence on $n$, since it is clear from context, and just write $\mathbf{X},\mathbf{y}$ and $\boldsymbol{\beta}$ with the indexing implicit. Furthermore, the following hold:
  1. A1
    lim $\limsup\lambda_{\max}(\mathbf{D}_{n})<C$ almost surely for some constant $C$.
  2. A2
    Each $\mathbf{X}_{n}$ has dimensions $n\times p(n)$, with $p(n)/n\xrightarrow{n\to\infty}\gamma\in[0,\infty)$.
  3. A3
    Each entry of $\boldsymbol{\epsilon}$ is independent with finite moment of order $4+\eta$, for some $\eta>0$. Furthermore, $\mathbb{E}[\epsilon_{i}]=0$ and $\mathbb{E}[\epsilon_{i}^{2}]=\sigma^{2}$.
  4. A4
    To fix scaling and simplify our results, we assume without loss of generality that $\frac{n^{\prime}\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{np}=1$.
  5. A5
    Each $\boldsymbol{\beta}_{n}$ is fixed or random independent of $\mathbf{X}_{n}$.
  6. A6
    $\widetilde{\mathbf{X}}$ is an independent draw from a potentially different right-rotationally invariant ensemble.
In all that follows, we state only the main results and defer proofs to the appendix.
3.1 An Analysis of Cross-Validation Methods
In [31], it is shown under a broad class of designs with i.i.d. rows that both LOOCV and GCV are asymptotically (under proportional asymptotics) consistent for the out-of-sample error. In our case, LOOCV is not tractable to analyze without additional assumptions on the distribution of $\mathbf{Q}$ – see Appendix B.2 for details. However, analysis of GCV, which was proposed as a rotationally invariant alternative to the standard leave-one-out cross-validation, is tractable. In particular, the GCV cross-validation metric is as follows:
$\mathsf{GCV}_{n}(\lambda)=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{y_{i}-\mathbf{x}_{i}^{\top}\hat{\boldsymbol{\beta}}_{\lambda}}{1-\mathsf{Tr}(\mathbf{S}_{\lambda})/n}\right)^{2}=\frac{1}{n-\mathsf{Tr}(\mathbf{S}_{\lambda})}\sum_{i=1}^{n}\left(y_{i}-\mathbf{x}_{i}^{\top}\hat{\boldsymbol{\beta}}_{\lambda}\right)^{2}.$ (4)
Here $\mathbf{S}_{\lambda}$ is the smoother matrix given by $\mathbf{X}^{\top}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}$. Our first result is as follows:
Theorem 1.
Under the stated assumptions,
$\mathsf{GCV}_{n}(\lambda)-\frac{r^{2}(v_{\mathbf{D}}(-\lambda)-\lambda v^{\prime}_{\mathbf{D}}(-\lambda))+\sigma^{2}\gamma v^{\prime}_{\mathbf{D}}(-\lambda)}{\gamma v_{\mathbf{D}}(-\lambda)^{2}}\xrightarrow{a.s.}0.$ (5)
We also find that for right-rotationally invariant designs, the asymptotic risk takes the following form:
Theorem 2.
Fix $0<\lambda_{1}<\lambda_{2}<\infty$. Under the stated assumptions,
$\sup_{\lambda\in[\lambda_{1},\lambda_{2}]}\left|R_{\mathbf{X},\mathbf{y}}\left(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta}\right)-\underbrace{\left(r^{2}\left(\frac{\lambda^{2}}{\gamma}v^{\prime}_{\mathbf{D}}(-\lambda)+\frac{\gamma-1}{\gamma}\right)+\sigma^{2}(v_{\mathbf{D}}(-\lambda)-\lambda v_{\mathbf{D}}^{\prime}(-\lambda))\right)}_{\mathsf{R}_{\mathbf{D}}(r^{2},\sigma^{2})}\right|\xrightarrow{a.s.}0.$ (6)
where we recall that $r^{2}=\|\boldsymbol{\beta}_{n}\|^{2}/n$ and $\sigma^{2}=\mathbb{E}{\epsilon_{i}^{2}}$.
The classical GCV then produces consistent tuning results if asymptotic value of GCV as a function of $\lambda$ (cf. Theorem 1) equals to the asymptotic risk as a function of $\lambda$ (cf. Theorem 2) up to some monotonic transform. This is what is observed for i.i.d. data under relatively lax assumptions, shown in [19] for the isotropic setting with random $\boldsymbol{\beta}$ and more generally in [31], but these proofs rely explicitly on identities given by the Silverstein equation [36, 19, 31]. Such identities do not exist outside of this i.i.d. setting, and the GCV is naturally biased once the data violates this assumption. Precisely, the asymptotic value of $\mathsf{GCV}_{n}$ in Theorem 1 is not necessarily equal to the asymptotic risk $\mathsf{R}_{\mathbf{D}}$ in Theorem 2 (up to an additive factor of $\sigma^{2}$) outside of the i.i.d. setting. We present an alternative approach which is provably consistent for right-rotationally invariant designs.
3.2 A Modified GCV
To obtain a consistent estimate of $R_{\mathbf{X},\mathbf{y}}$, our idea is to estimate $r^{2}$ and $\sigma^{2}$ and use these to compute the expression $\mathsf{R}_{\mathbf{D}}(r^{2},\sigma^{2})$ given by Theorem 2. We begin by discussing how estimate these two parameters.
3.2.1 Estimators for $r^{2},\sigma^{2}$
Estimating $\sigma^{2}$ in these settings has already been studied in [27], wherein a consistent estimator for $\sigma^{2}$ is presented. Our approach, which we observe to have better finite sample performance, will be slightly different. The original GCV can be seen as starting with the training error and finding a rescaling that recovers the out-of-sample risk. We likewise begin with the training error, but instead use it to produce estimating equations. In particular, we establish the following result:
Lemma 2.
Under the stated assumptions, for any $\lambda>0$,
$\frac{1}{n}\mathbf{y}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}\mathbf{y}-\left[r^{2}\frac{\lambda^{2}}{\gamma}\left(v_{\mathbf{D}}(-\lambda)-\lambda v^{\prime}_{\mathbf{D}}(-\lambda)\right)+\sigma^{2}\lambda^{2}v^{\prime}_{\mathbf{D}}(-\lambda)\right]\xrightarrow{a.s.}0.$ (7)
Observe that equation (7) can be used as an estimating equation, meaning that, if one treats  (7) as an equality to 0 (i.e. replacing “$\xrightarrow{a.s.}$” with “$=$”), then we can solve for the the only unknown quantities are $r^{2}$ and $\sigma^{2}$, allowing us to solve for them. The first term, $\mathbf{y}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}\mathbf{y}/n$ is the training error, then $\gamma=p/n$ along with $v_{\mathbf{D}}$ and $v^{\prime}_{\mathbf{D}}$, defined in Section 2.1, are all data dependent quantities. Moreover, it is an estimating equation for every value of $\lambda$, and each value yields a different estimating equation. While it takes only two values of $\lambda$ for us to solve for $r^{2}$ and $\sigma^{2}$, we find that this method is quite sensitive to the choice of $\lambda$. Instead, we choose a grid of $\lambda$ and find OLS estimates for $r^{2}$ and $\sigma^{2}$, which reduces this sensitivity. Explicitly, the scheme proceeds as shown at the start of Algorithm 1.
A simple corollary of Lemma 2 is then the following:
Corollary 3.
$\hat{r}^{2}$ and $\hat{\sigma}^{2}$ are strongly consistent for $r^{2}$ and $\sigma^{2}$, respectively.
It should be possible to show that these estimators have $O(\sqrt{n})$ rates under stronger assumptions on the distribution of the noise, such as sub-Gaussianity, but this analysis is not pursued.
Algorithm 1 ROTI-GCV($\lambda$)
1:$\mathbf{X},\mathbf{y}$, regularization strength $\lambda$; for use in estimation, grid of regularization strengths $\lambda_{1},\dots,\lambda_{L}$;
2: Estimate for ${R}_{\mathbf{X},\mathbf{y}}(\hat{\boldsymbol{\beta}},\boldsymbol{\beta})$
3:for $i=1,\dots,L$ do
4:     $a_{i}\leftarrow\frac{\lambda^{2}_{i}}{\gamma}\left(v_{\mathbf{D}}(-\lambda_{i})-\lambda_{i}v^{\prime}_{\mathbf{D}}(-\lambda_{i})\right)$
5:     $b_{i}\leftarrow\lambda_{i}^{2}v^{\prime}_{\mathbf{D}}(-\lambda_{i})$
6:     $t_{i}\leftarrow\|\mathbf{y}-\mathbf{X}\hat{\boldsymbol{\beta}}_{\lambda_{i}}\|_{2}^{2}.$
7:end for
8:$\hat{\sigma}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X},\mathbf{y})\leftarrow\frac{\sum_{i=1}^{L}\left(\frac{t_{i}}{a_{i}}-\frac{t_{1}}{a_{1}}\right)\left(\frac{b_{i}}{a_{i}}-\frac{b_{1}}{a_{1}}\right)}{\sum_{i=1}^{L}\left(\frac{b_{i}}{a_{i}}-\frac{b_{1}}{a_{1}}\right)^{2}}$,   $\hat{r}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X},\mathbf{y})\leftarrow\frac{\sum_{i=1}^{L}\left(\frac{t_{i}}{b_{i}}-\frac{t_{1}}{b_{1}}\right)\left(\frac{a_{i}}{b_{i}}-\frac{a_{1}}{b_{1}}\right)}{\sum_{i=1}^{L}\left(\frac{a_{i}}{b_{i}}-\frac{a_{1}}{b_{1}}\right)^{2}}.$
9:return $\mathsf{R}_{\mathbf{D}}(\hat{r}^{2},\hat{\sigma}^{2})$
3.2.2 Plug-in estimation
We now simply define $\mathsf{ROTI}\text{-}\mathsf{GCV}(\lambda)=\mathsf{R}_{\mathbf{D}}(\hat{r}^{2},\hat{\sigma}^{2})$ (recall $\mathsf{R}_{\mathbf{D}}$ from Theorem 2) and tune $\lambda$ using this metric; furthermore, this quantity serves as an estimator for the out-of-sample risk. It is then clear from Theorem 2 that this is consistent on compact intervals.
Corollary 4.
Fix $0<\lambda_{1}<\lambda_{2}<\infty$. Under Assumptions A1-A6, one has
$\sup_{\lambda\in[\lambda_{1},\lambda_{2}]}\left|\mathsf{ROTI}\text{-}\mathsf{GCV}(\lambda)-R_{\mathbf{X},\mathbf{y}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})\right|\xrightarrow{a.s.}0.$ (8)
Remark 1.
Replacing $\hat{r}^{2}$ and $\hat{\sigma}^{2}$ with any other set of consistent estimators also produces a valid estimate for the risk.
4 GCV under Signal-PC Alignment
An issue that emerges when attempting to apply GCV for right-rotationally invariant ensembles in practice is that the signal often tends to align with the top eigenvectors of the data, violating the independence assumption imposed on $\boldsymbol{\beta}$ and $\mathbf{O}$. The right-rotationally invariant assumption, for fixed $\boldsymbol{\beta}$, inherently assumes that the signal is incoherent with respect to the eigenbasis. However, this assumption is generally not true, and in fact in the i.i.d. anisotropic case, the geometry of $\boldsymbol{\beta}$ with respect to the top eigenvectors of $\boldsymbol{\Sigma}$ is known to significantly influence the behavior of ridge regression (see e.g. [45]).
Adapting the approach given in [27], we model this scenario using two index sets $\mathcal{J}_{a}$ and $\mathcal{J}_{c}$ of distinguished eigenvectors, which are aligned and coupled eigenvectors, respectively. We replace the last two assumptions in Section 3 with the following:
  1. A7
    The true signal $\boldsymbol{\beta}$ is given as $\boldsymbol{\beta}=\boldsymbol{\beta}^{\prime}+\sum_{i\in\mathcal{J}_{a}}\sqrt{n}\alpha_{i}\mathbf{o}_{i}$, where $\|\boldsymbol{\beta}\|/\sqrt{n}=r$. An eigenvector $\mathbf{o}_{i}$ is an aligned eigenvector if $\mathbf{o}_{i}\in\mathcal{J}_{a}$, as then $\mathbf{o}_{i}$ aligns with the signal $\boldsymbol{\beta}$.
  2. A8
    The test data $\widetilde{\mathbf{X}}=\widetilde{\mathbf{Q}}^{\top}\widetilde{\mathbf{D}}\widetilde{\mathbf{O}}$ is drawn from a separate right-rotationally invariant ensemble, where $\widetilde{\mathbf{O}}$ is distributed according to a Haar matrix where the eigenvectors are conditioned to satisfy $\mathbf{o}_{i}=\widetilde{\mathbf{o}}_{i}$ for all $i\in\mathcal{J}_{c}$. An eigenvector $\mathbf{o}_{i}$ is then coupled if $i\in\mathcal{J}_{c}$, as $\mathbf{o}_{i}$ is shared between the test and train sets.
We require two more assumptions:
  1. A9
    $\mathcal{J}_{a},\mathcal{J}_{c}$, and the $\boldsymbol{\alpha}=(\alpha_{i})_{i\in\mathcal{J}_{a}}$ are fixed and do not change with $n,p$.
  2. A10
    lim $\limsup\|\mathbb{E}[\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}}]\|_{\mathsf{op}}<C$ for some constant $C$.
The main difficulty that arises in our setting, compared to [27], is that we must account for how the geometry of the eigenvectors of the test set relates to those of the training set. One should expect, for example, in some types of structured data that the top eigenvectors of the training data and test data are closely aligned. We account for this effect through the coupled eigenvectors condition (Assumption A8). Furthermore, note that without this coupling, any alignment between the eigenvectors of the training set and the signal $\boldsymbol{\beta}$ will not exist in the test set, since then the eigenvectors of the test set would be independent of this alignment. We are still interested in the same test risk $R_{\mathbf{X},\mathbf{y}}$ as in 2, except now note that $\widetilde{\mathbf{X}}$ and $\mathbf{X}$ are not independent.
Theorem 3.
Under Assumptions A1-A4 and A7-A10, for any $0<\lambda_{1}<\lambda_{2}<\infty$, we have
$\sup_{\lambda\in[\lambda_{1},\lambda_{2}]}\left|R_{\mathbf{X},\mathbf{y}}\left(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta}\right)-\mathcal{R}_{\mathbf{X},\mathbf{y}}(r^{2},\sigma^{2},\boldsymbol{\alpha})\right|\xrightarrow{a.s.}0,$ (9)
where
$\mathcal{R}_{\mathbf{X},\mathbf{y}}(r^{2},\sigma^{2},\boldsymbol{\alpha})=\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})+\mathcal{V}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta}).$ (10)
Denote $\mathfrak{d}_{i}^{2}=\mathbb{E}\left[\left(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}}\right)_{ii}\right]$ and $\mathfrak{d}_{\mathsf{bulk}}^{2}=\frac{1}{p-|\mathcal{J}_{c}|}\sum_{i\not\in\mathcal{J}_{c}}\mathfrak{d}_{i}^{2}$. Then
$\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\lambda^{2}}{n}\sum_{i=1}^{p}\frac{\left(r^{2}/\gamma+\mathds{1}(i\in\mathcal{J}_{a})n\alpha_{i}^{2}\right)\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)}{(D_{ii}^{2}+\lambda)^{2}}$
$\mathcal{V}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\sigma^{2}}{n}\left[\sum_{i=1}^{n}\frac{D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)\right]$
Remark 2.
The assumption that the top eigenvectors of training and test sample covariance are exactly equal is unrealistic (Assumption A8) and is not expected to hold in practice. However, the idea is that the approximation can lead to more robust procedures. The covariance structure imposed on the test data under the assumption of coupled eigenvectors takes the form $\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\mathbf{o}_{i}\mathbf{o}_{i}^{\top}+\mathfrak{d}_{\mathsf{bulk}}^{2}\mathbf{I}$ (computed using [33, Lemma 4]). This can be interpreted as a low-rank approximation of the covariance of the test data with its bulk filled in with an identity piece. We thus expect this approximation to hold when the training and test data have “spiky” spectra, where the top few eigenvalues are heavily separated from the bulk, and these top eigenvectors of the test and training data are close.
4.1 Estimating $r^{2}$, $\sigma^{2}$, and $\alpha_{i}$
Algorithm 2 aROTI-GCV($\lambda$)
1:$\mathbf{X},\mathbf{y}$, regularization strength $\lambda$; index sets $\mathcal{J}_{a},\mathcal{J}_{c}$; consistent estimators $\hat{\mathfrak{d}}_{i}^{2}$ and $\hat{\mathfrak{d}}_{\mathsf{bulk}}^{2}$; for use in estimation of $r^{2},\sigma^{2}$, grid of regularization strengths $\lambda_{1},\dots,\lambda_{L}$;
2: Estimate for ${R}_{\mathbf{X},\mathbf{y}}(\hat{\boldsymbol{\beta}},\boldsymbol{\beta})$
3:$\mathbf{X}_{\mathsf{proj}}\leftarrow\mathbf{X}\mathbf{O}_{\mathcal{J}_{a}}^{\top}$
4:$\hat{\boldsymbol{\alpha}}\leftarrow(\mathbf{X}_{\mathsf{proj}}^{\top}\mathbf{X}_{\mathsf{proj}})^{-1}\mathbf{X}_{\mathsf{proj}}^{\top}\mathbf{y}$
5:$\mathbf{y}_{\mathsf{new}}=(\mathbf{D}_{\bar{\mathcal{J}_{a}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}_{a}}})^{1/2}(\mathbf{X}_{\mathsf{resid}}^{\top}\mathbf{X}_{\mathsf{resid}})^{-1}\mathbf{X}_{\mathsf{resid}}^{\top}\mathbf{y}$
6:$\mathbf{X}_{\mathsf{new}}=(\mathbf{D}_{\bar{\mathcal{J}_{a}}}^{\top}\mathbf{D}_{\bar{\mathcal{J}_{a}}})^{1/2}\mathbf{O}_{\bar{\mathcal{J}_{a}}}$
7:return $\mathcal{R}_{\mathbf{X},\mathbf{y}}\left(\hat{r}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}}),\hat{\sigma}^{2}_{\{\lambda_{i}\}_{i=1}^{L}}(\mathbf{X}_{\mathsf{new}},\mathbf{y}_{\mathsf{new}}),\hat{\boldsymbol{\alpha}},\{\hat{\mathfrak{d}}\}_{i\in\mathcal{J}_{c},\mathsf{bulk}}\right)$
Our approach is simply to estimate the $\alpha_{i}$’s using principle components regression (see [25, 23, 4, 22, 10, 2, 21, 47, 24, 35, 37] for details on principle components regression), described in steps 1 and 2 below. We then compute a reduced model which is right-rotationally invariant, from which vanilla $\mathsf{ROTI}\text{-}\mathsf{GCV}$ can be used to find $r^{2}$ and $\sigma^{2}$.
Throughout all that follows, we focus on subsets $S$ of the indices of all nonzero singular values $\mathcal{N}=\{i\in[p]:(\mathbf{D}^{\top}\mathbf{D})_{ii}>0\}$. Let $\mathbf{O}_{S}\in\mathbb{R}^{|S|\times p}$ for any set $S$ denote the rows of $\mathbf{O}$ indexed by $S$. Let $\mathbf{D}_{S}\in\mathbb{R}^{(n-|S|)\times p}$ denote the rows of $\mathbf{D}$ indexed by $S$. Finally, for a set $S\subset\mathcal{N}$, denote $\overline{S}=\mathcal{N}\setminus S$. The procedure then proceeds as in steps 1-4 of Algorithm 2.
Lemma 5.
Under the additional assumption that $\boldsymbol{\epsilon}$ has Gaussian entries, as described above, $\hat{r}^{2}$, $\hat{\sigma}^{2}$, and $\hat{\boldsymbol{\alpha}}$ are all strongly consistent.
The assumption of Gaussian entries is technical and we believe it can be removed with some work.
4.2 Cross-validation metric
Our approach to forming a cross-validation metric is to then again explicitly compute the risk formulas
$\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\lambda^{2}}{n}\sum_{i=1}^{p}\frac{\left(r^{2}/\gamma+\mathds{1}(i\in\mathcal{J}_{a})n\alpha_{i}^{2}\right)\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)}{(D_{ii}^{2}+\lambda)^{2}},$
$\mathcal{V}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\sigma^{2}}{n}\left[\sum_{i=1}^{n}\frac{D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)\right],$
where we estimate the unknown parameters from data. To apply our method, we therefore need to consistently estimate $r^{2},\sigma^{2},\boldsymbol{\alpha}$, and $\{\mathfrak{d}_{i}^{2}\}_{i\in\mathcal{J}_{c}}$ and $\mathfrak{d}_{\mathsf{bulk}}^{2}$. Lemma 5 gives us most of these parameters, but we still require consistent estimates of the $\mathfrak{d}_{i}$’s. This is a nonissue in cases where $\mathbf{D}$ and $\widetilde{\mathbf{D}}$ are modeled as deterministic. If they are random, we expect that if the test data is drawn from the same right-rotationally invariant ensemble as the training data, then, in practice, one can use the eigenvalues of the training data to do this (for theoretical guarantees under some examples of random $\mathbf{D}$, see results on eigenvalue rigidity [9]). However, our framework further allows for the test distribution to be drawn from a different right-rotationally invariant ensemble. In such cases, it may be possible to then use test data to do this estimation because it only depends on the test covariates $\widetilde{\mathbf{X}}$ (which the practitioner may have access to), but not its labels. We therefore make one final assumption:
  1. A11
    We assume there exists estimators $\hat{\mathfrak{d}}_{i}$ and $\hat{\mathfrak{d}}_{\mathsf{bulk}}$ where we have $\sup_{i\in\mathcal{J}_{a}\cup\mathcal{J}_{c}}|\hat{\mathfrak{d}}_{i}-\mathfrak{d}_{i}|\xrightarrow{a.s.}0$, and $|\hat{\mathfrak{d}}_{\mathsf{bulk}}-\mathfrak{d}_{\mathsf{bulk}}|\xrightarrow{a.s.}0$.
We then take our validation metric as $\mathsf{aROTI}\text{-}\mathsf{GCV}(\lambda)=\mathcal{R}_{\mathbf{X},\mathbf{y}}(\hat{r}^{2},\hat{\sigma}^{2},\hat{\boldsymbol{\alpha}},\{\hat{\mathfrak{d}}_{i}\}_{i\in\mathcal{J}_{c},\mathsf{bulk}})$, where the latter expression denotes substituting every parameter with its estimated version.
Corollary 6.
Under Assumptions A1-A4, A7-A11, and the additional assumption that the entries of $\boldsymbol{\epsilon}$ are Gaussian, $\mathsf{aROTI}\text{-}\mathsf{GCV}$ satisfies
$\sup_{\lambda\in[\lambda_{1},\lambda_{2}]}|\mathsf{aROTI}\text{-}\mathsf{GCV}(\lambda)-R_{\mathbf{X},\mathbf{y}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})|\xrightarrow{a.s.}0.$ (11)
The total algorithm was described above in Algorithm 2.
On selecting $\mathcal{J}_{a},\mathcal{J}_{c}$:
[27] presents a way to do hypothesis testing for alignment to detect $\mathcal{J}_{a}$. For detecting $\mathcal{J}_{c}$, one can measure the angle among the top eigenvectors of the test set and top eigenvectors of the training set, placing pairs of vectors which have overlap deviating heavily from $90^{\circ}$ into $\mathcal{J}_{c}$.
5 Numerical Experiments
In this section, we demonstrate the efficacy of our method via numerical experiments. In producing each plot below, we generate a training set according to a prescribed distribution, as well as a signal vector $\boldsymbol{\beta}$. We then repeatedly resample the training noise $\boldsymbol{\epsilon}$ and run each cross-validation method, in addition to computing the true risk on the test set.
5.1 Performance of ROTI-GCV on right-rotationally invariant data
Figure 1 illustrates the performance of ROTI-GCV and compares it to LOOCV and GCV. In these figures, tuned risk refers to the out-of-sample prediction error obtained when we use the value of $\lambda$ that optimizes the given cross-validation method. For the Expected MSE, we report the minimum of the true expected out-of-sample mean squared error curve. Standard errors are shown for all tuned risks. Est Risk refers to the value of the out-of-sample risk that would be estimated using the given cross-validation method. As expected, in the Gaussian setting all three methods perform well. However, for dependent data, such as in the autocorrelated and equicorrelated cases, we observe that the risk curves produced by GCV and LOOCV are wildly off from the truth, whereas ROTI-GCV accurately captures the true risk curve. Surprisingly, the tuned risk values from LOOCV, GCV are still decently close to that of ROTI-GCV. Because these two methods estimate the total mean squared error $\mathbb{E}[(y-\mathbf{x}^{\top}\hat{\boldsymbol{\beta}})^{2}\mid\mathbf{X},\mathbf{y}]=R_{\mathbf{X}}(\hat{\boldsymbol{\beta}},\boldsymbol{\beta})+\sigma^{2}$ , the values they produce include an additional factor of the noise level $\sigma^{2}$, while the risk we are interested (See equation (2)) does not include this term, and hence are plotted with this term removed.
(a) Gaussian data
(b) Autocorrelated data, $\rho=0.9$
(c) $t$-distributed data
(d) Equicorrelated data, $\rho=0.8$
Figure 1: Risk curves produced by the cross-validation methods (LOOCV, GCV, ROTI-GCV), in addition to the true risk curve. Each plot uses a different form of right-rotationally invariant data, as indicated in each caption. (b): Autocorrelated data: rows of $\mathbf{X}$ are drawn according to $\mathbf{x}_{i}=\rho\mathbf{x}_{i-1}+\sqrt{1-\rho^{2}}\boldsymbol{\epsilon}_{i}$, with $\boldsymbol{\epsilon}_{i}$ being i.i.d. draws from $\mathsf{N}(0,\mathbf{I}_{n})$. We set $\rho=0.8$. (c): $t$-distributed data: each row is drawn from a multivariate $t$ distribution with $3$ degrees of freedom. (d): Equicorrelated data $\mathbf{X}\in\mathbb{R}^{n\times p}$ has independent columns, but each column follows a multivariate Gaussian distribution with covariance matrix $\boldsymbol{\Sigma}$, where $\Sigma_{ij}=\rho$ if $i\neq j$, and $\Sigma_{ii}=1$. All simulations have $n=p=1000$ and $r^{2}=\sigma^{2}=1$. The $x$-axis for every plot is $\lambda$, the regularization parameter. The colored lines (blue, green, red, purple) are one of 10 iterations. In each iteration, we compute the cross-validation metric over a range of $\lambda$ to produce the line, which reflects the estimated out-of-sample risk. Theoretical risk is plotted in all three as reference (black line), and refers to the exact value of $R_{\mathbf{X}}(\hat{\boldsymbol{\beta}},\boldsymbol{\beta})$ (see equation 2). The dashed blue line shows the average expected MSE curve as a benchmark.
5.2 Experiments under Signal-PC alignment
As discussed in Section 4, the scenario where the signal $\boldsymbol{\beta}$ aligns with the eigenvectors $\mathbf{O}$ is problematic for right-rotationally invariant designs. In Figure 2, we show that the aROTI-GCV method from Section 4 works when our assumptions are met in a somewhat contrived example where we set the top eigenvectors of the test set to be identical to those of the training set, matching our coupled eigenvectors condition. In Figure 3, we give some examples in which this situation can arise.
Figure 2 illustrates the accuracy of our method when this alignment is introduced. Here, the signal aligns with the top 10 eigenvectors, and these top 10 eigenvectors are also coupled with the test set. As expected, the theory matches when these assumptions are met.
Figure 2: GCV under Signal-PC alignment. As in Figure 1, curves show the risk and estimated risk as a function of $\lambda$. $\mathbf{X}$ is drawn with autocorrelated rows. We set $\rho=0.8$. The top 10 eigenvectors ($\mathcal{J}_{c}=[10]$) are coupled, so that the top 10 eigenvectors of the test set are equal to those of $\mathbf{X}$. $\boldsymbol{\beta}$ is drawn according to $\boldsymbol{\beta}^{\prime}+\sqrt{n}\sum_{i=1}^{10}\frac{i}{10}\mathbf{o}_{i}$, with $\|\boldsymbol{\beta}^{\prime}\|^{2}=1$, $\sigma^{2}=1$, $n=p=1000$.
[…]
(a) i.i.d. rows drawn from a Gaussian mixture, i.e. $\mathbf{x}_{i}$ is drawn from $\frac{1}{2}\mathsf{N}(\mathbf{3},\mathbf{I}_{p})+\frac{1}{2}\mathsf{N}(-\mathbf{3},\mathbf{I}_{p})$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=\{1\}$.
(b) i.i.d. rows drawn from an equicorrelated Gaussian, i.e. $\mathbf{x}_{i}$ is drawn from $\mathsf{N}(0,\boldsymbol{\Sigma})$, where $\Sigma_{ij}=\mathds{1}(i=j)+\rho\mathds{1}(i\neq j)$. We set $\rho=0.5$. When computing ROTI-GCV, we set $\mathcal{J}_{c}=[10]$.
Figure 3: GCV performance on data with correlation. As in Figure 1, curves show the risk and estimated risk as a function of $\lambda$.
6 Discussion and Conclusion
Takeaways: We introduce ROTI-GCV, a method for estimating out-of-sample prediction error under a new model for data, namely right-rotationally invariant covariates. The model departs crucially from the existing framework of having i.i.d. samples drawn from a given distribution and sub-Gaussian tailed covariates. Instead it allows for structured forms of dependence and heavy tails. ROTI-GCV is a new cross-validation framework that is provably consistent for these designs and we demonstrate its superiority over existing methods across diverse settings. As a byproduct of our analysis, we produce new estimators for signal-to-noise ratio and noise variance that are robust to these deviations from the i.i.d. assumptions. At the core, our idea involves identifying the relation given by Lemma 2 and noticing that the term on the right involves $\lambda$. This means we can iterate and compute this term over a grid of values of $\lambda$. This leads to a series of estimating equations from which stable estimators of $r^{2},\sigma^{2}$ can be derived. This core idea can be applicable for choosing hyperparameters for much broader models beyond linear regression–pretty much anywhere an analogue of Lemma 2 is available. As future steps, extending our method to other settings such as generalized linear models, kernel ridge regression, regularized random features regression would be important directions. Furthermore, recent advances in [8, 44] show that right-rotationally invariant distributions have a broad universality class–in fact as long as the singular vectors of a design matrix satisfy certain generic structure, properties of learning algorithms under such settings are well captured by results derived under the right-rotationally invariant assumption. In particular, [8, 44] establish such universality properties of the in-sample risk for ridge regression. We expect analogous universality results to hold in our setting. Thus, ROTI-GCV should apply broadly beyond right-rotationally invariant designs. In fact, it should be effective for any design that belongs to the spectral universality class identified by [8]. A comprehensive exploration of how this universality can be used to improve hyperparameter tuning is a promising direction for future research.
Limitations: Right-rotationally invariant designs may on occasion have difficulty modeling distributions with feature covariances that are much more complicated than those covered in Section 5, especially when the spectrum is not dominated by a few top eigenvectors. Finding ways to account for different types of feature covariance in data using these designs would naturally increase the broader applicability of our method.
7 Acknowledgements
P.S. would like to acknowledge partial funding from NSF DMS-2113426.
Appendix A Deferred Proofs
A.1 Proof of Lemma 1
Proof.
Let $\mathbf{O}$ have rows $\mathbf{o}_{i}$. Then
$\mathbb{E}[\mathbf{X}^{\top}\mathbf{X}]$ $=\mathbb{E}[\mathbf{O}^{\top}\mathbf{D}^{\top}\mathbf{D}\mathbf{O}]=\sum_{i=1}^{n\wedge p}\mathbb{E}[D_{ii}^{2}\mathbf{o}_{i}\mathbf{o}_{i}^{\top}]=\sum_{i=1}^{n\wedge p}\mathbb{E}[D_{ii}^{2}]\mathbb{E}[\mathbf{o}_{i}\mathbf{o}_{i}^{\top}]$
=$\stackrel{{\scriptstyle(*)}}{{=}}\frac{1}{p}\mathbf{I}_{p}\sum_{i=1}^{n\wedge p}\mathbb{E}[D_{ii}^{2}]=\mathsf{Tr}(\mathbb{E}[\mathbf{D}^{\top}\mathbf{D}])/p\cdot\mathbf{I}_{p}$
where ()(*)( ∗ ) follows from $\mathbf{I}_{p}=\mathbb{E}[\mathbf{O}^{\top}\mathbf{O}]=\sum_{i=1}^{p}\mathbb{E}[\mathbf{o}_{i}\mathbf{o}_{i}^{\top}]$ and the exchangeability of the $\mathbf{o}_{i}$ (which is derived from the fact that $\mathbb{O}(p)$ contains the permutation matrices). ∎
A.2 Useful Results
Lemma 7 (Hanson-Wright for Spherical Vectors).
Let $\mathbf{v}\in\mathbb{R}^{n}$ be a random vector distributed uniformly on $S^{n-1}$ and let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be a fixed matrix. Then
$\mathbb{P}(\left|\mathbf{v}^{\top}\mathbf{A}\mathbf{v}-\mathbb{E}[\mathbf{v}^{\top}\mathbf{A}\mathbf{v}]\right|\geq t)\leq 2\exp\left(-C\min\left(\frac{n^{2}t^{2}}{2K^{4}\|\mathbf{A}\|_{F}^{2}},\frac{nt}{K^{2}\|\mathbf{A}\|}\right)\right).$
for some absolute constants $C,K$.
Proof.
Isoperimetric inequalities imply that vectors uniformly distributed on $S^{n-1}$ satisfy convex concentration (see [34]) with constant $K/\sqrt{n}$, for some absolute constant $K$, where we note that having sub-Gaussian tails around the median is equivalent to having them around the mean (see e.g. [5]). We can then apply [1, Theorem 2.3] and rescale. ∎
Lemma 8 (Expectations of Quadratic Forms).
Let $\mathbf{v}$ be uniformly distributed on $S^{n-1}$, and let $\mathbf{A}$ be an $n\times n$ matrix. Then
$\mathbb{E}[\mathbf{v}^{\top}\mathbf{A}\mathbf{v}]=\mathsf{Tr}(\mathbf{A}\mathbb{E}[\mathbf{v}\mathbf{v}^{\top}])=\frac{\mathsf{Tr}(\mathbf{A})}{n}$ (12)
Proof.
The first equality is clear. The second can be seen to follow from the observation used in the proof of Lemma 1 and the fact that columns of a Haar distributed matrix are uniform on the sphere (see e.g. [30]). Alternatively, recall that for $\mathbf{g}\sim\mathsf{N}(0,\mathbf{I}_{n})$, $\frac{\mathbf{g}}{\|\mathbf{g}\|}\perp\!\!\!\perp\|\mathbf{g}\|$. Hence
$\mathbb{E}[\mathbf{v}\mathbf{v}^{\top}]=\mathbb{E}[\mathbf{g}\mathbf{g}^{\top}/\|\mathbf{g}\|^{2}]=\mathbb{E}[\mathbf{g}\mathbf{g}^{\top}]/\mathbb{E}[\|\mathbf{g}\|^{2}]=\mathbf{I}_{n}/n.$
Lemma 9.
Let $\mathbf{v}_{n}$ be a sequence of random vectors, where $\mathbf{v}_{n}$ is uniformly distributed on $S^{n-1}$. Let $\mathbf{u}_{n}$ be another sequence of random vectors independent of $\mathbf{v}_{n}$, where lim $\limsup\|\mathbf{u}_{n}\|<C$ for some constant $C$ almost surely. Then $\mathbf{v}_{n}^{\top}\mathbf{u}_{n}\xrightarrow{a.s.}0$.
Proof.
We replace $\mathbf{v}_{n}$ with $\frac{\mathbf{g}}{\|\mathbf{g}\|}$ where $\mathbf{g}\sim\mathsf{N}(0,\mathbf{I}_{p}/p)$. Then conditional on $\mathbf{u}_{n}$, we have $\mathbf{g}^{\top}\mathbf{u}_{n}\sim\mathsf{N}(0,\|\mathbf{u}_{n}\|^{2}/p)$. Some standard tail bounds complete this.
$\mathbb{P}\left(|\mathbf{g^{\top}\mathbf{T}_{4}}|>t\mid\mathbf{u}_{n}\right)$ $\leq\sqrt{\frac{2}{\pi}}\frac{\|\mathbf{u}_{n}\|}{t\cdot\sqrt{p}}\exp\left(-t^{2}/2\cdot\frac{p}{\|\mathbf{u}_{n}\|^{2}}\right).$
Since $\limsup_{n\to\infty}\|\mathbf{u}_{n}\|$ is a.s. bounded, the bound above is summable in $n$, so by Borel-Cantelli, the convergence is almost sure. An application of Slutsky finishes, since we know $\|\mathbf{g}\|\xrightarrow{a.s.}1$ by the strong Law of Large Numbers. ∎
A.3 Proof of Theorem 1
We analyze the numerator and denominator separately. The latter is simple:
$(1-\mathsf{Tr}(\mathbf{S}_{\lambda})/n)^{2}$ $=\left(1-\mathsf{Tr}((\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})/n\right)^{2}=\left(\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda}{\lambda+D_{ii}^{2}}\right)^{2}=\gamma^{2}\lambda^{2}v_{\mathbf{D}}(-\lambda)^{2}$
The numerator only requires slightly more work. Recall $\mathbf{y}=\mathbf{X}\boldsymbol{\beta}+\boldsymbol{\epsilon}$. Furthermore, note that the GCV numerator can be rewritten as below:
$\frac{1}{n}\mathbf{y}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}\mathbf{y}=\underbrace{\frac{1}{n}\boldsymbol{\beta}^{\top}\mathbf{X}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}\mathbf{X}\boldsymbol{\beta}}_{T_{1}}+2\underbrace{\frac{1}{n}\boldsymbol{\epsilon}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}\mathbf{X}\boldsymbol{\beta}}_{T_{2}}+\underbrace{\frac{1}{n}\boldsymbol{\epsilon}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}\boldsymbol{\epsilon}}_{T_{3}}.$
We handle each term. First, $T_{1}$ simplifies easily into a quadratic form which allows us to apply Hanson-Wright (Lemma 7):
$T_{1}$ $=\frac{1}{n}(\mathbf{O}\boldsymbol{\beta})^{\top}\mathbf{D}^{\top}(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})^{2}\mathbf{D}(\mathbf{O}\boldsymbol{\beta})$
This again concentrates around its expectation, which is
$=\frac{\|\boldsymbol{\beta}\|^{2}}{n}\frac{\mathsf{Tr}(\mathbf{D}^{\top}(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})^{2}\mathbf{D})}{p}$
$=\frac{\|\boldsymbol{\beta}\|^{2}}{n}\frac{1}{p}\sum_{i=1}^{p}D_{ii}^{2}\left(1-\frac{D_{ii}^{2}}{D_{ii}^{2}+\lambda}\right)^{2}$
$=\frac{\|\boldsymbol{\beta}\|^{2}}{n}\frac{\lambda^{2}}{p}\sum_{i=1}^{p}\frac{D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}$
$=\frac{\|\boldsymbol{\beta}\|^{2}}{n}\lambda^{2}\left(m_{\mathbf{D}}(-\lambda)-\lambda m^{\prime}_{\mathbf{D}}(-\lambda)\right)=\frac{\|\boldsymbol{\beta}\|^{2}}{n}\frac{\lambda^{2}}{\gamma}\left(v_{\mathbf{D}}(-\lambda)-\lambda v^{\prime}_{\mathbf{D}}(-\lambda)\right)$
Letting $\mathbf{P}=\mathbf{D}^{\top}(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})^{2}\mathbf{D}$, Hanson-Wright (Lemma 7)implies
$\mathbb{P}(|T_{1}-\mathbb{E}[T_{1}]|>t)\leq 2\exp\left(-c\min\left(\frac{p^{2}t^{2}}{2K^{4}\|\mathbf{P}\|_{F}^{2}},\frac{pt}{K^{2}\|\mathbf{P}\|}\right)\right)=2\exp\left(p\min\left(\frac{t^{2}}{2K^{4}\frac{1}{p}\|\mathbf{P}\|_{F}^{2}},\frac{t}{K^{2}\|\mathbf{P}\|}\right)\right)$
Note that
$\mathbf{P}=\mathrm{diag}\left(\frac{\lambda^{2}D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}\right)_{i=1}^{p}$
Hence
$\frac{1}{p}\|\mathbf{P}\|_{F}^{2}$ $=\frac{1}{p}\sum_{i=1}^{p}\left(\frac{\lambda^{2}D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}\right)^{2}=\frac{1}{p}\sum_{i=1}^{p}\left(\frac{\lambda^{2}}{(D_{ii}^{2}+\lambda)^{2}}\right)^{2}D_{ii}^{4}\leq\|\mathbf{D}\|_{\mathsf{op}}^{4}$
$\|\mathbf{P}\|_{\mathsf{op}}$ $=\frac{\lambda^{2}D_{11}^{2}}{(D_{11}^{2}+\lambda)^{2}}\leq\|\mathbf{D}\|^{2}$
By assumption, both terms are almost surely bounded in the limit. Recalling that $p(n)/n\to\gamma$, the above bound is summable in $n$ and hence Borel-Cantelli implies almost sure convergence.
To handle the second term, we use a slightly different method.
$T_{2}$ $=\frac{1}{n}{\boldsymbol{\epsilon}}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}\mathbf{X}\boldsymbol{\beta}$
$=\frac{1}{n}(\mathbf{O}\boldsymbol{\beta})^{\top}\underbrace{\mathbf{D}^{\top}\mathbf{Q}(\mathbf{I}-\mathbf{S}_{\lambda})^{2}{\boldsymbol{\epsilon}}}_{\mathbf{T}_{4}}$
Note that $\tilde{\boldsymbol{\beta}}=\mathbf{O}\boldsymbol{\beta}$ is a uniformly random vector of norm $\|\boldsymbol{\beta}\|$. We replace this with $\|\boldsymbol{\beta}\|\frac{\mathbf{g}}{\|\mathbf{g}\|}$ where $\mathbf{g}\sim\mathsf{N}(0,\mathbf{I}_{p}/p)$. Then conditional on $\mathbf{D},\boldsymbol{\epsilon}$, we have $\frac{1}{\sqrt{n}}\mathbf{g}^{\top}\mathbf{T}_{4}\sim\mathsf{N}(0,\|\mathbf{T}_{4}\|/(pn))$. Furthermore note $\|\mathbf{T}_{4}\|\leq\|\mathbf{D}^{\top}\|\|\mathbf{Q}\|\|\mathbf{I}-\mathbf{S}_{\lambda}\|^{2}\|\boldsymbol{\epsilon}\|\leq\|\mathbf{D}^{\top}\|\|\boldsymbol{\epsilon}\|$. Some standard tail bounds complete this.
$\frac{1}{n}\|\boldsymbol{\beta}\|\frac{\mathbf{g}^{\top}}{\|\mathbf{g}\|}\mathbf{T}_{4}$ $=\frac{\|\boldsymbol{\beta}\|}{\sqrt{n}}\frac{1}{\|\mathbf{g}\|}\left(\frac{1}{\sqrt{n}}\mathbf{g}^{\top}\mathbf{T}_{4}\right)=r\cdot\frac{1}{\|\mathbf{g}\|}\frac{1}{\sqrt{n}}\mathbf{g^{\top}\mathbf{T}_{4}}$
Hence
$\mathbb{P}\left(|\frac{1}{\sqrt{n}}\mathbf{g^{\top}\mathbf{T}_{4}}|>t\mid\mathbf{D},\boldsymbol{\epsilon}\right)$ $\leq\sqrt{\frac{2}{\pi}}\frac{\|\mathbf{T}_{4}\|}{t\cdot\sqrt{pn}}\exp\left(-t^{2}/2\cdot\frac{pn}{\|\mathbf{T}_{4}\|^{2}}\right).$
We claim that $\limsup_{n\to\infty}\|\mathbf{T}_{4}\|/\sqrt{n}$ is bounded. This follows from the bound above, the law of large numbers applied to $\boldsymbol{\epsilon}$, and the assumption on $\mathbf{D}$. Hence the bound above is summable in $n$ once more, so by Borel-Cantelli, the convergence is almost sure. An application of Slutsky finishes, since we know $\|\mathbf{g}\|\xrightarrow{a.s.}1$ by the strong Law of Large Numbers. The convergence for $T_{3}$ follows directly from [6, Lemma C.3]
A.4 Proof of Theorem 2
Proof of Theorem 2.
In order to understand how one should tune the regularization parameter in this problem, we first must understand the out of sample risk for a given value of $\lambda$. The risk admits a decomposition into three terms:
$\frac{1}{n^{\prime}}\mathbb{E}\left[\left\|\widetilde{\mathbf{X}}(\boldsymbol{\beta}-\hat{\boldsymbol{\beta}})\right\|_{2}^{2}\mid\mathbf{X},\mathbf{y}\right]$ $=\frac{1}{n^{\prime}}\cdot\frac{\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{p}\left\|\boldsymbol{\beta}-\hat{\boldsymbol{\beta}}\right\|_{2}^{2}$
$=\frac{\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{n^{\prime}p}\|(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})\boldsymbol{\beta}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\boldsymbol{\epsilon}\|_{2}^{2}$
$=\frac{n\mathbb{E}[\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})]}{n^{\prime}p}\left[\frac{1}{n}\boldsymbol{\beta}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})^{2}\boldsymbol{\beta}\right.$
$\qquad\left.+\frac{1}{n}\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-2}\mathbf{X}^{\top}\boldsymbol{\epsilon}-2\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-2}\mathbf{X}^{\top}\mathbf{X}\boldsymbol{\beta}\right]$
The normalization assumption insures that the prefactor is $1$. The first term corresponds to bias, the second to the variance, and the third is negligible. For the bias,
$B_{\mathbf{X}}(\hat{\boldsymbol{\beta}},\boldsymbol{\beta})$ $=n^{-1}\boldsymbol{\beta}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})^{2}\boldsymbol{\beta}$
$=n^{-1}(\mathbf{O}\boldsymbol{\beta})^{\top}\left[\mathbf{I}-(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D}\right]^{2}(\mathbf{O}\boldsymbol{\beta})$
$=n^{-1}(\mathbf{O}\boldsymbol{\beta})^{\top}\left(\lambda(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\right)^{2}(\mathbf{O}\boldsymbol{\beta})$
Again let $\mathbf{b}=(\mathbf{O}\boldsymbol{\beta})/\|\boldsymbol{\beta}\|$ and $\mathbf{P}=\left(\lambda(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\right)^{2}$, and note $\mathsf{Tr}(\mathbf{P})=\sum_{i=1}^{p}\frac{\lambda^{2}}{(D_{ii}^{2}+\lambda)^{2}}$ and $\|\mathbf{P}\|_{F}^{2}=\sum_{i=1}^{p}\frac{\lambda^{4}}{(D_{ii}^{2}+\lambda)^{4}}$. Furthermore, one has $\mathbb{E}[\mathbf{b}^{\top}\mathbf{P}\mathbf{b}]=\frac{1}{p}\mathsf{Tr}(\mathbf{P})$ using 8, and clearly $\|\mathbf{P}\|\leq 1$. We now apply Lemma 7 to show concentration:
$\mathbb{P}\left(\left|\mathbf{b}^{\top}\mathbf{P}\mathbf{b}-\frac{1}{p}\mathsf{Tr}(\mathbf{P})\right|\geq t\right)$ $\leq 2\exp\left(-C\min\left(\frac{p^{2}t^{2}}{2K^{4}\|\mathbf{P}\|^{2}_{F}},\frac{pt}{K^{2}}\right)\right)$
$=2\exp\left(-Cp\min\left(\frac{t^{2}}{2K^{4}\tfrac{1}{p}\|\mathbf{P}\|^{2}_{F}},\frac{t}{K^{2}}\right)\right)$
Note that $\frac{1}{p}\|\mathbf{P}\|^{2}_{F}$ is always bounded above by $1$, so this bound can never be made vacuous by a certain setting of $\mathbf{D}$. This bound is summable in $n$ and thus we have almost sure convergence of the difference to zero.
This yields pointwise convergence for the bias. To strengthen this to uniform convergence, we show that the derivative of the difference $\mathbf{b}^{\top}\mathbf{P}\mathbf{b}-\frac{1}{p}\mathsf{Tr}(\mathbf{P})$ is almost surely bounded on $[\lambda_{1},\lambda_{2}]$:
$\frac{\mathrm{d}}{\mathrm{d}\lambda}\mathbf{b}^{\top}\mathbf{Pb}=\frac{\mathrm{d}}{\mathrm{d}\lambda}\left[\frac{1}{n}\sum_{i=1}^{p}(\mathbf{o}_{i}^{\top}\boldsymbol{\beta})^{2}\frac{\lambda^{2}}{(D_{ii}^{2}+\lambda)^{2}}\right]$ $=\frac{1}{n}\sum_{i=1}^{p}(\mathbf{o}_{i}^{\top}\boldsymbol{\beta})^{2}\frac{2\lambda(D_{ii}^{2}+\lambda)^{2}+2(D_{ii}^{2}+\lambda)(\lambda^{2})}{(D_{ii}^{2}+\lambda)^{4}}$
$\leq\frac{1}{n}\frac{2\lambda(D_{11}^{2}+\lambda)^{2}+2(D_{11}^{2}+\lambda)(\lambda^{2})}{(\lambda)^{4}}\sum_{i=1}^{p}(\mathbf{o}_{i}^{\top}\boldsymbol{\beta})^{2}$
$=\frac{1}{n}\frac{2\lambda(D_{11}^{2}+\lambda)^{2}+2(D_{11}^{2}+\lambda)(\lambda^{2})}{(\lambda)^{4}}$
which is indeed almost surely bounded on $[\lambda_{1},\lambda_{2}]$ in the limit by assumption on $\|\mathbf{D}\|$. Similarly, $\frac{\mathrm{d}}{\mathrm{d}\lambda}(\frac{1}{p}\mathsf{Tr}(\mathbf{P}))$ is also bounded on this interval by the same quantity, meaning the difference is Lipschitz. Hence one can discretize the interval $[\lambda_{1},\lambda_{2}]$ into a sufficiently fine grid. Pointwise convergence on every point holds at every point on the grid, then the Lipschitz condition assures that the convergence is uniform.
For the third term, we rewrite it as
$2(\boldsymbol{\epsilon}/\sqrt{n})^{\top}\mathbf{Q}\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-2}\mathbf{D}^{\top}\mathbf{D}\cdot(\mathbf{O}\boldsymbol{\beta}/\sqrt{n})$
and then apply Lemma 9 to see that this converges almost surely to zero. We can likewise control the derivative uniformly over the interval and again obtain almost-sure convergence.
Lastly, for the variance, we first compute
$\mathbb{E}\left[\frac{1}{n}\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-2}\mathbf{X}^{\top}\boldsymbol{\epsilon}\right]=\frac{\sigma^{2}}{n}\sum_{i=1}^{n\wedge p}\frac{D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}.$ (13)
Then by [6, Lemma C.3], this term satisfies
$\mathbb{E}\left[\frac{1}{n}\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-2}\mathbf{X}^{\top}\boldsymbol{\epsilon}\right]-\frac{1}{n}\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-2}\mathbf{X}^{\top}\boldsymbol{\epsilon}\xrightarrow{a.s.}0.$ (14)
The derivative is again uniformly controlled, producing uniform convergence once more.
A.5 Proof of Corollary 4
Proof.
Under the assumptions of Theorem 1, for any consistent estimators $(\hat{r}^{2},\hat{\sigma}^{2})$ of $(r^{2},\sigma^{2})$, respectively, one has, for any $\lambda_{0}>0$,
[…]
If $\hat{r}^{2},\hat{\sigma}^{2}$ are strongly consistent, then naturally the convergence is almost sure. This is immediate from the fact that
$\lambda^{2}v^{\prime}_{\mathbf{D}}(-\lambda)=\frac{1}{n}\sum_{i=1}^{n}\frac{\lambda^{2}}{(D_{ii}^{2}+\lambda)^{2}}$ $\leq 1$
$v_{\mathbf{D}}(-\lambda)-\lambda v^{\prime}_{\mathbf{D}}(-\lambda)=\frac{1}{n}\sum_{i=1}^{n\wedge p}\frac{D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}$ $\leq\frac{1}{4\lambda}.$
Combining Equation (15) with Theorem 2 then completes the corollary. ∎
A.6 Proof of Lemma 2
This was proven in the course of Theorem 1, as the training error is nothing but the numerator of the original GCV.
A.7 Proof of Corollary 3
This follows from the fact that the method is continuous in the $t_{i}$ and we have pointwise convergence from Lemma 2.
A.8 Proof of Theorem 3
Proof.
Most computations are analogous to the proof of Theorem 2. We first compute $\mathbb{E}[\widetilde{\mathbf{X}}^{\top}\widetilde{\mathbf{X}}\mid\mathbf{X},\mathbf{y}]$ using [33, Lemma 4]. We find that it is equal to $\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2})\mathbf{o}_{i}\mathbf{o}_{i}^{\top}+\mathfrak{d}_{\mathsf{bulk}}^{2}\mathbf{I}$. Call this quantity $\mathbf{C}$ for convenience. We then move to computing the risk.
$\boldsymbol{\beta}-\hat{\boldsymbol{\beta}}=(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})\boldsymbol{\beta}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\boldsymbol{\epsilon}.$ (16)
Meanwhile
$\mathbb{E}\left[\|\widetilde{\mathbf{X}}(\boldsymbol{\beta}-\hat{\boldsymbol{\beta}})\|_{2}^{2}\right]=(\boldsymbol{\beta}-\hat{\boldsymbol{\beta}})^{\top}\left(\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2})\mathbf{o}_{i}\mathbf{o}_{i}^{\top}+\mathfrak{d}_{\mathsf{bulk}}^{2}\mathbf{I}\right)(\boldsymbol{\beta}-\hat{\boldsymbol{\beta}}).$ (17)
Thus as before, there are three main terms to handle.
Variance term:
$\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{C}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\boldsymbol{\epsilon}$
$=\mathfrak{d}_{\mathsf{bulk}}^{2}\boldsymbol{\epsilon}^{\top}\mathbf{Q}^{\top}\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-2}\mathbf{D}^{\top}\mathbf{Q}\boldsymbol{\epsilon}+$
$\qquad\boldsymbol{\epsilon}^{\top}\mathbf{Q}^{\top}\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\left(\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{ii}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\mathbf{e}_{i}\mathbf{e}_{i}^{\top}\right)(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{Q}\boldsymbol{\epsilon}.$
Note now that
$\frac{n}{n^{\prime}}\mathfrak{d}_{\mathsf{bulk}}^{2}=\frac{n}{n^{\prime}}\frac{1}{p-|\mathcal{J}_{c}|}\sum_{i\not\in\mathcal{J}_{c}}\mathfrak{d}_{i}^{2}=\frac{n\mathbb{E}\mathsf{Tr}(\widetilde{\mathbf{D}}^{\top}\widetilde{\mathbf{D}})}{n^{\prime}p}+O(1/p)=1+O(1/p),$ (18)
where the penultimate equality is due to assumption A10 and the last is due to the scaling assumption A4. Note that the $O(1/p)$ term converges almost surely to zero.
Then
$\frac{n}{n^{\prime}}\mathfrak{d}_{\mathsf{bulk}}^{2}\left|\frac{1}{n}\boldsymbol{\epsilon}^{\top}\mathbf{Q}^{\top}\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-2}\mathbf{D}^{\top}\mathbf{Q}\boldsymbol{\epsilon}-\frac{1}{n}\mathbb{E}[\boldsymbol{\epsilon}^{\top}\mathbf{Q}^{\top}\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-2}\mathbf{D}^{\top}\mathbf{Q}\boldsymbol{\epsilon}]\right|$
$\leq\frac{}{}|\boldsymbol{\epsilon}^{\top}\mathbf{Q}^{\top}\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-2}\mathbf{D}^{\top}\mathbf{Q}\boldsymbol{\epsilon}-\mathbb{E}[\boldsymbol{\epsilon}^{\top}\mathbf{Q}^{\top}\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-2}\mathbf{D}^{\top}\mathbf{Q}\boldsymbol{\epsilon}]|+O(1/p)\xrightarrow{a.s.}0.$
The same arguments as in Theorem 2 then imply uniform convergence to zero.
On the other hand, the second term is order $1$ and hence vanishes when divided by $n$. The convergence is then uniform because the function is monotonic in $\lambda$. It is almost sure through, for instance, using [6, Lemma C.3]. Note that the form of $\mathcal{V}_{\mathbf{X}}$ contains an extra
$\frac{\sigma^{2}}{n}\sum_{i\in\mathcal{J}_{c}}\frac{D_{ii}^{2}}{D_{ii^{2}}+\lambda}\mathfrak{d}_{i}^{2}$
term which is uniformly negligible by monotonicity and the assumption that $\mathfrak{d}_{i}^{2}$ is a.s. bounded in the limit.
Bias term:
$\boldsymbol{\beta}^{\top}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})^{\top}\mathbf{C}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X}))\boldsymbol{\beta}$
$=\mathfrak{d}_{\mathsf{bulk}}^{2}(\mathbf{O}\boldsymbol{\beta})^{\top}(\mathbf{I}-(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})^{2}(\mathbf{O}\boldsymbol{\beta})$
$\qquad+(\mathbf{O}\boldsymbol{\beta})^{\top}\left[(\mathbf{I}-(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1})\left[\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2})\mathbf{e}_{i}\mathbf{e}_{i}^{\top}\right](\mathbf{I}-(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\right](\mathbf{O}\boldsymbol{\beta}).$
We now need to handle this a bit more carefully than before, because $\boldsymbol{\beta}$ is no longer independent of $\mathbf{O}$. Note that $\mathbf{O}\boldsymbol{\beta}=\mathbf{O}\boldsymbol{\beta}^{\prime}+\sum_{i\in\mathcal{J}_{a}}\alpha_{i}\mathbf{e}_{i}$. There are a few different types of terms we need to handle.
  1. 1.
    Unaligned terms:
    $\mathfrak{d}_{\mathsf{bulk}}^{2}(\mathbf{O}\boldsymbol{\beta}^{\prime})^{\top}(\mathbf{I}-(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})^{2}(\mathbf{O}\boldsymbol{\beta}^{\prime})$
    $\qquad+(\mathbf{O}\boldsymbol{\beta})^{\top}\left[(\mathbf{I}-(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})\left[\sum_{i\in\mathcal{J}_{c}}(\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2})\mathbf{e}_{i}\mathbf{e}_{i}^{\top}\right](\mathbf{I}-(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})\right](\mathbf{O}\boldsymbol{\beta})$
    The first term is treated equivalently to the bias in the ridge case in Theorem 2, where, as in the variance computation above, the $\mathfrak{d}_{\mathsf{bulk}}^{2}$ can be handled. This produces uniform convergence to its expectation. The second term is $O(|\mathfrak{J}_{c}|)$ and hence vanishes when divided by $n$. It does so uniformly because it is monotone increasing in $\lambda$.
  2. 2.
    Aligned but uncoupled terms:
    $\sum_{i\in\mathcal{J}_{a}\setminus\mathcal{J}_{c}}n\alpha_{i}^{2}\mathfrak{d}_{\mathsf{bulk}}^{2}\left(1-\frac{D_{ii}^{2}}{D_{ii}^{2}+\lambda}\right)$
  3. 3.
    Aligned and coupled terms:
    $\sum_{i\in\mathcal{J}_{a}\cap\mathcal{J}_{c}}n\alpha_{i}^{2}\mathfrak{d}_{i}^{2}\left(1-\frac{D_{ii}^{2}}{D_{ii}^{2}+\lambda}\right)$
  4. 4.
    Cross terms:
    $\sum_{i\in\mathcal{J}_{a}}\mathbf{o}_{i}^{\top}\boldsymbol{\beta}^{\prime}\sqrt{n}\alpha_{i}\left[(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{ii}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2}))\left(1-\frac{D_{ii}^{2}}{D_{ii}^{2}+\lambda}\right)\right]$
    These terms are only of order $O(\sqrt{n})$ and hence vanish - they do so uniformly because, after taking absolute value of the summand, the resulting term is monotone increasing in $\lambda$, and thus it suffices to control the value at $\lambda_{2}$.
    Note that the cross terms between $\alpha_{i}\mathbf{e}_{i}$ and $\alpha_{j}\mathbf{e}_{j}$ are all identically zero and thus do not need to be considered.
Summing all of these together, one obtains uniform convergence to $\mathcal{B}$. Note that the statement of $\mathcal{B}$ contains an extra
$\frac{\lambda^{2}}{n}\sum_{i\in\mathcal{J}_{c}}\frac{r^{2}}{\gamma}\frac{\mathfrak{d}_{i}^{2}-\mathfrak{d}_{\mathsf{bulk}}^{2}}{(D_{ii}^{2}+\lambda)^{2}}$ (19)
term, which is uniformly negligible.
Cross term:
$\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{C}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})\boldsymbol{\beta}$
$=\boldsymbol{\epsilon}^{\top}\mathbf{X}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{C}(\mathbf{I}-(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{\top}\mathbf{X})(\boldsymbol{\beta}^{\prime}+\sum_{i\in\mathcal{J}_{a}}\alpha_{i}\mathbf{o}_{i})$
$=\boldsymbol{\epsilon}^{\top}\mathbf{Q}\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{O}\mathbf{C}\mathbf{O}^{\top}(\mathbf{I}-\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{D})\mathbf{O}\boldsymbol{\beta}^{\prime}+$
$\qquad\sum_{i\in\mathcal{J}_{a}\cap\mathcal{J}_{c}}\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}\frac{D_{ii}}{\lambda+D_{ii}^{2}}\cdot\mathfrak{d}_{i}^{2}\cdot\frac{\lambda}{D_{ii}^{2}+\lambda}\cdot\sqrt{n}\alpha_{i}$
For the first term, note that $\mathbf{O}\mathbf{C}\mathbf{O}^{\top}$ is in fact independent of $\mathbf{O}$. Moreover, the matrix in the middle has bounded operator norm, and thus this term vanishes after dividing by $1/n$ due to Lemma 9, where we apply the strong law of large numbers to bound $\|\boldsymbol{\epsilon}\|$. For the second, we note that this is almost surely, for sufficiently large $n$, uniformly bounded above by
$C\sqrt{n}\cdot(\max_{i\in\mathcal{J}_{a}}\alpha_{i})\cdot\max_{i\in\mathcal{J}_{a}\cap\mathcal{J}_{c}}|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|$
Recall we are dividing everything by $n$. Hence if we show that $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|/\sqrt{n}\xrightarrow{a.s.}0$, then we will be done with a union bound. We rewrite $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|=\boldsymbol{\epsilon}^{\top}(\mathbf{q}_{i}\mathbf{q}_{i}^{\top})\boldsymbol{\epsilon}$ and can now apply [6, Lemma C.3] to find
$\frac{1}{n}\boldsymbol{\epsilon}^{\top}(\mathbf{q}_{i}\mathbf{q}_{i}^{\top})\boldsymbol{\epsilon}-\frac{\sigma^{2}}{n}\xrightarrow{a.s.}0.$
Thus $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|^{2}/n$ converges a.s. to 00, and thus by continuous mapping so does $|\mathbf{q}_{i}^{\top}\boldsymbol{\epsilon}|/\sqrt{n}$.
A.9 Proof of Lemma 5
The following is taken from [27]. We note that in this work, the errors are assumed Gaussian. We will mark which ones require this and how we believe they can be relaxed. First, the consistency of $\alpha_{i}$ is immediate from the first result of [27, (40)], where we note that the proof does not require Gaussianity of the errors. The one step in the proof where care must be made is in controlling
$\frac{1}{p}\left\|\mathbf{O}_{\mathcal{J}}^{\top}\left(\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{D}_{\mathcal{J}}\right)^{-1}\mathbf{D}_{\mathcal{J}}^{\top}\mathbf{Q}\varepsilon\right\|_{2}^{2},$ (20)
but we note that this can be done through [6, Lemma C.3].
Then the reduced model is right-rotationally invariant by [27, p. 58, (124)]. Showing that the reduced model is right-rotationally invariant requires Gaussianity and is what leads to the added assumption of Lemma 5. Without the Gaussian assumption, the errors in the reduced model will not be i.i.d. However, in general we only require that certain inner products concentrate. We believe this will still hold in the reduced model when one explicitly tracks the form of the error.
A.10 Proof of Corollary 6
We will prove uniform convergence term by term. We begin with the variance. Define the following:
$A$ $=\frac{1}{n}\left[\sum_{i=1}^{n}\frac{D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)\right]$
$\hat{A}$ $=\frac{1}{n}\left[\sum_{i=1}^{n}\frac{D_{ii}^{2}}{(D_{ii}^{2}+\lambda)^{2}}\left(\hat{\mathfrak{d}}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\hat{\mathfrak{d}}_{i}^{2}-\hat{\mathfrak{d}}^{2}_{\mathsf{bulk}})\right)\right]$
$T_{1}$ $=\sigma^{2}A\qquad T_{2}$ $=\hat{\sigma}^{2}\hat{A}\qquad T_{3}$ $=\hat{\sigma}^{2}{A}$
Note that $T_{1}=\mathcal{V}_{\mathbf{X}}$. First, note that $\hat{A}$ and $\hat{\sigma}^{2}/n$ are both almost surely bounded for sufficiently large $n$. This follows from the fact that the estimators are strongly consistent for bounded quantities. Note that the fact that we allow $o(n)$ fluctuations on $\hat{\mathfrak{d}}_{i}$ is allowed because there are finitely many of these, and we later divide by $n$. We can thus uniformly control $T_{1}-T_{3}$ and then $T_{3}-T_{2}$, producing uniform convergence.
For the bias, we wish to uniformly control
$\mathcal{B}_{\mathbf{X}}(\hat{\boldsymbol{\beta}}_{\lambda},\boldsymbol{\beta})$ $=\frac{\lambda^{2}}{n}\sum_{i=1}^{p}\frac{\left(r^{2}/\gamma+\mathds{1}(i\in\mathcal{J}_{a})n\alpha_{i}^{2}\right)\left(\mathfrak{d}_{\mathsf{bulk}}^{2}+\mathds{1}(i\in\mathcal{J}_{c})(\mathfrak{d}_{i}^{2}-\mathfrak{d}^{2}_{\mathsf{bulk}})\right)}{(D_{ii}^{2}+\lambda)^{2}}.$
One can then expand the numerator into 4 constituent terms. Arguments similar to that of the variance then show uniform convergence for each of them.
A.11 Weak Convergence of Spectral Measure
Remark 3.
In cases such as the i.i.d. case, the diagonal entries of $\mathbf{D}$ converge weakly almost surely to some limiting spectral distribution. Under the following assumption, one can, instead of stating results in terms of the difference in the predicted and actual risks converging almost surely to zero, as in Corollary 4, instead write them as converging almost surely to the Stieltjes transforms taken over the limiting measure:
  • (A*)
    Almost surely, the diagonal entries of $\mathbf{D}$ converge in the Wasserstein-2 ($W_{2}$) distance to a measure $\mathsf{D}$, where $\mathsf{D}^{2}$ has non-zero mean and compact support $\mathrm{supp}(\mathsf{D}^{2})\subseteq[0,\infty)$. See Appendix B.3 for details on Wasserstein-2 convergence.
Lemma 10.
Under the assumptions of Theorem 1 and Assumption (A*), then one has, for any $\lambda_{1},\lambda_{2}$ satisfying $0<\lambda_{1}<\lambda_{2}$ and any two consistent estimators $\hat{r}^{2},\hat{\sigma}^{2}$,
[…]
As usual, if the two estimators are instead strongly consistent, then the above convergence is almost sure.
This is then similar in flavor to [31, Theorem 4.1].
Appendix B Miscellaneous Details
B.1 Experimental details
All experiments were run on a personal laptop. Each experiment took only on the order of 1 minute. All experiments are carried out with $n=p=1000$, because this scenario is most sensitive to the amount of regularization needed (as with $\lambda=0$, often times the risk is infinite). In Section 3, we first sample a training set. We then repeatedly resample the training noise $\boldsymbol{\epsilon}$ because this is more economical. Each time, we then run each cross-validation technique and plot the loss curve over $\lambda$. The estimated risk is then the risk reported by the loss curve at its minimum. The tuned risk is then the actual mean-squared error (which has an analytical expression) produced when using the estimator with that value of $\lambda$. For ROTI-GCV, Assumption A4, while useful theoretically, proves to provide some practical challenges, mainly that finding the correct rescaling of the data that makes this assumption hold is not always simple, since the trace does not necessarily concentrate. In the experiments in Section 3, the data is drawn according to distributions satisfying this assumption, and the data is not normalized after being sampled. That said, normalizing the data does not affect any results. It is important to note that this scaling never affects the tuned value of $\lambda$ and instead only affects the risk estimate.
In Section 4, it becomes necessary to estimate the eigenvalues of the test set data. To do this, we now sample a test set ($n=p=1000$ again), and then compute the eigenvalues of the test set. A similar result should hold when using eigenvalues of the training set, since here both sets are equal in distribution.
B.2 LOOCV Intractability
The LOOCV in our setting is difficult to analyze because the objective depends on $\mathbf{Q}$, on which we do not place any assumptions. In particular, one can derive
$\mathsf{LOOCV}_{n}(\lambda)$ $=\mathbf{y}^{\top}(\mathbf{I}-\mathbf{S}_{\lambda})\mathbf{D}_{\lambda}^{-2}(\mathbf{I}-\mathbf{S}_{\lambda})\mathbf{y}/n$
$=\mathbf{y}^{\top}(\mathbf{I}-\mathbf{Q}^{\top}(\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})\mathbf{Q})\mathbf{D}_{\lambda}^{-2}(\mathbf{I}-\mathbf{Q}^{\top}(\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})\mathbf{Q})\mathbf{y}/n$
where $\mathbf{D}_{\lambda}=\mathrm{diag}((1-(\mathbf{S}_{\lambda})_{ii})_{i=1}^{n})$. We try to simplify the diagonal term first.
$(D_{\lambda})_{ii}$ $=1-\mathbf{x}_{i}^{\top}(\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{x}_{i}$
$=1-\mathbf{q}_{i}^{\top}\mathbf{D}\mathbf{O}(\mathbf{O}^{\top}\mathbf{D}^{\top}\mathbf{D}\mathbf{O}+\lambda\mathbf{I})^{-1}\mathbf{O}^{\top}\mathbf{D}^{\top}\mathbf{q}_{i}$
$=1-\mathbf{q}_{i}^{\top}\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top}\mathbf{q}_{i}$
where $\mathbf{q}_{i}$ is the $i$-th row of $\mathbf{Q}$. We analyze
$\mathbf{Q}\mathbf{D}^{-2}_{\lambda}\mathbf{Q}^{\top}$ $=\sum_{i=1}^{n}\mathbf{q}_{i}\mathbf{q}_{i}^{\top}(\mathbf{D}_{\lambda}^{-2})_{ii}.$
Then $\mathsf{LOOCV}$ should contribute two terms, as the cross term should vanish. They are
$(\mathbf{O}\boldsymbol{\beta})^{\top}\mathbf{D}^{\top}\mathbf{Q}(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})(\mathbf{Q}\mathbf{D}_{\lambda}^{-2}\mathbf{Q}^{\top})(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})\mathbf{O}\boldsymbol{\beta}$
which should concentrate around its trace.
The second is
$\boldsymbol{\epsilon}^{\top}(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\mathbf{D}^{\top})(\mathbf{Q}\mathbf{D}_{\lambda}^{-2}\mathbf{Q}^{\top})(\mathbf{I}-\mathbf{D}(\mathbf{D}^{\top}\mathbf{D}+\lambda\mathbf{I})^{-1}\boldsymbol{\epsilon}$
which also concentrates around some multiple of its trace.
Finding this trace now involves numerous terms with $\mathbf{Q}$. Everything except the center term is diagonal, but immediately one arrives at the following difficulty:
$(\mathbf{Q}\mathbf{D}^{-2}_{\lambda}\mathbf{Q})_{ii}$ $=\sum_{j=1}^{n}(\mathbf{q}_{j}\mathbf{q}_{j}^{\top})_{ii}(\mathbf{D}_{\lambda}^{-2})_{jj}$
$=\sum_{j=1}^{n}(\mathbf{q}_{ji})^{2}(\mathbf{D}_{\lambda}^{-2})_{jj}$
Assuming that the rows of $\mathbf{Q}$ are exchangeable conditional on $\mathbf{D}$ is insufficient to make progress. If $\mathbf{Q}$ is assumed to also be Haar, then likely closed forms can be derived, but this is not pursued.
B.3 Wasserstein-2 Convergence
The following section is taken from [27, Definition 2.2]. For a matrix $(\mathbf{v}_{1},\dots,\mathbf{v}_{k})=(v_{i,1},\dots,v_{i,k})_{i=1}^{n}\in\mathbb{R}^{n\times k}$ and a random vector $(\mathsf{V}_{1},\dots,\mathsf{V}_{k})$, we write
$(\mathbf{v}_{1},\dots,\mathbf{v}_{k})\xrightarrow{W_{2}}(\mathsf{V}1_{,}\dots,\mathsf{V}_{k})$
to mean that the empirical distribution of the columns of $\mathbf{v}_{1},\dots,\mathbf{v}_{k})$ converge to $\mathsf{V}_{1},\dots,\mathsf{V}_{k}$ in Wasserstein-2 distance. This means that for any continuous function $f:\mathbb{R}^{k}\to\mathbb{R}$ satisfying
$|f(x,\dots,x_{k})|\leq C\left(1+\|(x_{1},\dots,x_{k})\|^{2}\right)$ (21)
for some $C>0$, we have
$\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}f(v_{i,1},\dots,v_{i,k})=\mathbb{E}\left[f(\mathsf{V}_{1},\dots,\mathsf{V}_{k})\right]$
where $\mathbb{E}\left[\|(\mathsf{V}_{1},\dots,\mathsf{V}_{k})\|^{2}\right]<\infty$.